OpenStudy (anonymous):

Use the demoivre's theorem to find the answer? (1+i)^8

OpenStudy (anonymous):

Let \(z=1+i\) then: \[\left| z \right|=\sqrt{1^2+1^2}=\sqrt 2; \theta= \tan^{-1}({1 \over 1})=\pi/4\]. Writing z in polar coordinates: \[z=\sqrt2 (\cos({\pi \over 4})+i \sin({\pi \over 4}))\]. Now apply demoivre's theorem: \[(1+i)^8=z^8=(\sqrt{2})^8(\cos(8{\pi \over 4})+i \sin(8{\pi \over 4}))=2^4(\cos(2\pi)+i \sin(2\pi))=16(1+i (0))=16\]

OpenStudy (anonymous):

By the way, you can easily prove this by simple calculation: \[(1+i)^8=[(1+i)^2]^4=(1+2i+i^2)^4=(2i)^4=2^4(i^4)=16\] Notice here that \(i^2=-1\) and \(i^4=1\).

OpenStudy (anonymous):

what if the question is (1-i)^11

OpenStudy (anonymous):

Try to follow the same procedure and tell what you get.

OpenStudy (anonymous):

what to do after this step 2(under root )^11 (cos-11/4+isin-11/4)

OpenStudy (anonymous):

sorry cos-11pie/4+isin-11pie/4

OpenStudy (anonymous):

\(cos (-11\pi/4)=sin(-11 \pi/4)=1/\sqrt2\)

OpenStudy (anonymous):

sorry \(-1/\sqrt2\)

OpenStudy (anonymous):

but how???

OpenStudy (anonymous):

\[\cos({-11 \pi \over 4})=\cos(4\pi-{11\pi \over 4})\cos({5\pi \over 4})\]. \(5\pi/4\) is the correspondent angle of \(pi/4\) in the third quadrant, Hence cos\((-11\pi/4)=-1/\sqrt2\).

OpenStudy (anonymous):

There is a missing equality in the first line between the two cosines :)